OpenStudy (icalibear):

How do you simplify (-2r^2s)^3 (3rs^2)?

OpenStudy (anonymous):

Example (-2 r^5 s^3)^2 = (-2)^2 r^10 s^ 6 = 4 r^10 s^6

OpenStudy (icalibear):

So you multiply the exponents?

OpenStudy (anonymous):

Another example ( a^2 b^3 c^7)^4= a^8 b^12 c^28

OpenStudy (icalibear):

by the -2

OpenStudy (icalibear):

So the answer to mine would be

OpenStudy (anonymous):

u add exponents when multiplying but when ( )^3 u multiply

OpenStudy (anonymous):

Try to do it

OpenStudy (icalibear):

oh okay

OpenStudy (icalibear):

okay (-2)^3 = 8 r^2 + 3 = r^5 s^1 + ^3 = s^4

OpenStudy (icalibear):

right?

OpenStudy (icalibear):

or i mean -8

OpenStudy (anonymous):

Your final answer should look like (-2 r^2 s)^3 (3 r s^2) = -24 r^7 s^5

OpenStudy (icalibear):

wait how did you get -24? sorry i just dont understand eachstep, when to multiply when to add etc

OpenStudy (anonymous):

3(-8)=-24

OpenStudy (anonymous):

(-2 r^2 s)^3=-8 r^6 s^3

OpenStudy (anonymous):

(-8 r^6 s^3) (3 r s^2) = -24 r^7 s^5

OpenStudy (anonymous):

Have a good day